3 Questions:

1)What is the pH of A solution containing 0.483 M potassium nitrite and 5.38E-2 M nitrous acid.

2)In the laboratory, a general chemistry student measured the pH of a 0.589 M aqueous solution of hypochlorous acid to be 3.861. Use the information she obtained to determine the Ka for this acid.

3)What is the pH of an aqueous solution of 0.421 M morphine
Answer #1

1) pH = pKa + log(salt/base)

         = 3.25 + log(0.483/0.0538)

         = 3.25 + 0.95

         = 4.2

2) [H+]= (Ka*C)^1/2     ([H+]= 1.37 x10^-4 from pH value given )

Ka *C = [H+]^1/2

          = (1.37 X10^-4)^1/2

Ka = (1.17 X10^-2)/0.589 = 1.98 X10^-2

3) [OH-] = (Kb*c)^1/2

             = (1.62 X 10^-6 *0.421)^1/2

             = 8.17 X 10^-4

pOH = -log(8.17 x10^-4)

        = 3.09

pH = 14-3.09 = 10.91